OpenStudy (anonymous):

What is the molarity of solution containing 5.035grams of FeCl3 in enough water to make 500ml of solution?

OpenStudy (anonymous):

molarity=no.of moles/volume in litre molartiy=(w/m)/v =w/(m*v) w=5.035gm; m=molecular wight of fecl3; v=500/1000 jst put the values nd u will get the answer

OpenStudy (doc.brown):

\[FeCl_3 = 55.845 + 3(35.45) = 162.20g/mol\]\[\frac{5.035gFeCl_3}{500mL}\times\frac{mol}{162.20g}\times\frac{1000mL}{L}=\]